how much money must be deposited today to become $1250 in 50 years at 4.5% compounded continuously?
Answer by stanbon(75887)

how much money must be deposited today to become $1250 in 50 years at 4.5% compounded continuously?
----------------
A(t) = Pe^(rt)
1250 = P*e^(0.045*50)
P = 1250/(e^0.045*50)
P = $131.75
=====================
Cheers,
Stan H.
==================